Homogenous Number System and Its Application in Physics

Mushfiq Ahmad University of Rajshahi, Bangladesh

In traditional number system 0 has a privileged position. It is the neutral number of addition, and it is the lone number, which does not have a reciprocal. We have developed a number system in which any arbitrarily chosen number can be the neutral number, and every number, including 0 has a reciprocal. This number system permits us to describe motion in terms of slowness – reciprocal velocity – just as it is possible to describe in terms of velocity. Reciprocal of 0 permits us to represent the slowness of the body at rest. Reciprocal representation permits us to state Einstein’s postulate in the form of Planck’s hypothesis. Homogenous number system also makes it possible to develop discrete energy levels from a classical treatment of simple harmonic oscillator. Thus Homogenous Number System provides a bridge between classical and quantum physics.